Circular No. 7945

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 O5 (NEAT) The following orbital elements from MPEC 2002-P06 are very uncer-

tain:

T = 2002 Aug. 2.873 TT  ω = 15.^{o}130
e = 0.64137  Ω = 281.707 }2000.0
q = 1.18274 AU  i = 21.102
a = 3.29795 AU n^{o} = 0.164566 P = 5.99 years

V4741 SAGITTARII S. Mazuk, C. C. Venturini, R. J. Rudy, D. K. Lynch, The Aerospace

Corporation; R. C. Puetter, Center for Astrophysics and Space Sciences, University of California at San Diego; and R. B. Perry, NASA Langley Research Center, report 0.8–2.5 μm spectroscopy of the nova V4741 Sgr using the Lick Observatory 3-meter telescope and Aerospace Corporation Near-Infrared Imaging Spectrograph on July 19.29 UT (∼ 95 days after discovery): “He I at 1.0830 μm is the strongest emission line in the spectrum and the H I Paschen and Brackett series are prominent. O I lines excited by Ly-β fluorescence are still present although He II lines are starting to appear. The emission lines are broad (FWHM ranging from 2600–3600 km/s) and have peculiar, flat topped profiles. No coronal lines are detected. The relative strengths of the He I and H I lines suggest that helium abundance is enhanced relative to the solar value.”

SUPERNOVA 2002ef IN NGC 7761 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002ef (cf. IAUC 7943), obtained by M. Calkins on Aug. 1.43 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a Type Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 7080 km/s for the host galaxy, the supernova expansion velocity is about 11 700 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 1 ± 2 days before maximum light.

2002 August 2 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother